Determine the pH of a solution consisting of 0.15 M NH4Cl and 1.5 M NH3. The Kb of NH3 is 1.8x10-5
.
1
Expert's answer
2020-12-01T14:23:26-0500

pH= ?

[NH4Cl]= 0.15M

[NH3]= 1.5M

Kb= 1.8 x 10-5

Now,

pKb= -logKb

pKb= -log(1.8x10-5)

pKb= 4.745

We can now calculate pOH using Henderson Hasselbalch equation

pOH= pKb + log[salt]/[Acid]

pOH= 4.745 + log(0.15/1.5)

pOH= 3.745

We can now find the pH

pH + pOH= 14

pH= 14 - pOH

pH= 14 - 3.745

pH= 10.26